OpenStudy (anonymous):

3 ( m - 6 ) + 2 ( m + 5 ) = 2; m =

OpenStudy (anonymous):

\[ 3 ( m - 6 ) + 2 ( m + 5 ) = 2 \rightarrow 3m - 18 + 2m + 10 = 2 \rightarrow 5m - 8 = 2 \rightarrow 5m = 10 \rightarrow m = \boxed {2}. \]

OpenStudy (anonymous):